The Avis Company is a car rental company and is located three miles from the Los Angeles airport (LAX). Avis is dispatching a bus from its offices to the airport every 2 minutes. The average traveling time (a round trip) is 20 minutes.

   

a.

How many Avis buses are traveling to and from the airport? (Round your answer to the nearest whole number.)

b.

The branch manager wants to improve the service and suggests dispatching buses every 0.5 minutes. She argues that this will reduce the average traveling time (a round trip) to 2.5 minutes . Is she correct? If your answer is negative, then what will the average traveling time be?
Answer #1

a.

The throughput rate is 30 buses per hour.

The throughput time is 20/60 = 1/3 hours

TT=WIP/TR

TT*TR=WIP

30*1/3=10

10 buses are traveling to and from the airport.


b.

This reduces the average waiting time, but has no effect on the average traveling time. The average traveling time does not change and will remain 20 min.
Answer #2

This reduces the average waiting time, but has no effect on the average traveling time. The average traveling time does not change and will remain 20 min
Answer #3

a> 20/2 = 10

but at time 20 mts avi had reached the airport so 9 bus


b> it remain same as distance remain same 20 mts